### v1

# Classical and quantum chaotic synchronization in coupled dissipative time crystals

Abstract
We investigate the dynamics of two coherently coupled dissipative time crystals. In the classical mean-field limit of infinite spin length, we identify a regime of chaotic synchronization, marked by a positive largest Lyapunov exponent and a Pearson correlation coefficient close to one. At the boundary of this regime, the Pearson coefficient varies abruptly, marking a crossover between staggered and uniform $z$-magnetization. To address finite-size quantum dynamics, we employ a quantum-trajectory approach and study the trajectory-resolved expectations of subsystem $z$-magnetizations. Their histograms over time and trajectory realizations exhibit maxima that undergo a staggered-to-uniform crossover analogous to the classical one. In analogy with the classical case, we interpret this behavior as quantum chaotic synchronization, with dissipative quantum chaos evidenced by the steady-state density matrix exhibiting Gaussian Unitary Ensemble statistics. The locations of the classical and quantum crossover differ, reflecting the noncommutativity of the infinite-spin and infinite-time limits and the impact of entanglement, quantified via the entanglement entropy between subsystems.
1 Introduction
Synchronization of weakly coupled nonlinear self-sustained oscillators is a ubiquitous phenomenon in nature, observed in systems ranging from coupled neon tubes, to heart cells, to people applauding in a theater [1]. An especially important case is the synchronization of many oscillators with slightly different frequencies, elegantly captured by the Kuramoto model [2, 3]. Recent progress in experimental control has brought quantum fluctuations to the forefront. Since the pioneering work of [4], quantum synchronization has been extensively studied in a variety of bipartite open quantum systems [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], as well as in many-body settings where collective quantum synchronization emerges [17, 18, 19, 20, 21], even showing robustness of Kuramoto synchronization to quantum noise [22]. Experimental realizations have demonstrated entrainment to external drives using cold atoms [23], trapped ions [24, 25], nuclear spins [26], and superconducting circuits [27]. In the classical domain, synchronization can also occur between chaotic oscillators [28, 29, 1, 30, 31], a phenomenon with no established quantum counterpart.
A related concept is that of time crystals, where a many-body system exhibits persistent oscillations in an order parameter, thus breaking continuous time-translation symmetry [32, 33]. In driven, unitary systems, this behavior appears as a collective subharmonic response [34, 35], while in dissipative systems it emerges as spontaneous oscillations of collective magnetization in the thermodynamic limit [36]. In this limit, such behavior often reduces to mean-field nonlinear dynamics with self-sustained oscillations. This naturally raises the question of whether coupled time crystals can synchronize. Following the first investigation of synchronization in many dissipative time crystals [37], subsequent works refined this picture [38], culminating in a recent study of synchronization between two dissipative (or boundary) time crystals in the thermodynamic limit [39]. These results revealed distinct synchronized phases, with oscillations that can be either in phase or out of phase.
In this contribution we work with the same model, focusing on signatures of chaos and synchronization in both classical and quantum regimes. In the mean-field infinite-spin limit, we characterize dynamics using the largest Lyapunov exponent (LLE) and the Pearson coefficient computed over long times and random initial conditions. We identify the onset of chaotic synchronization: as the coupling strength increases, the Pearson coefficient jumps toward a positive value near one in correspondence with the LLE becoming positive, indicating simultaneous onset of chaos and enhancement of synchronization, akin to chaotic synchronization of Lorenz systems [30, 31]. Alongside, the time-averaged $z$-magnetizations exhibit a crossover from staggered to uniform values across the two subsystems.
To probe quantum effects at finite spin magnitude, we employ the quantum-trajectory method and analyze trajectory-resolved $z$-magnetizations. Evaluating for each subsystem the histogram of the $z$-magnetization expectations, as they change over time and trajectory realizations, one can see a maximum that displays a sharp crossover from a staggered to a uniform regime across the two subsystems. This behavior closely echoes the classical case and suggests the emergence of quantum chaotic synchronization. This interpretation is strengthened by the steady-state density matrix, whose spectrum exhibits Gaussian unitary ensemble statistics that marks quantum chaos, and by the behavior of the bipartite entanglement entropy, highlighting the central role of entanglement in quantum synchronization [13, 4, 5, 40]. Notably, the crossover points in the quantum and classical cases do not coincide, reflecting the noncommutativity of the infinite-time and infinite-spin limits: While the classical system exhibits persistent oscillations, the quantum dynamics relaxes to a nonequilibrium steady state [41]. Nevertheless, both regimes display qualitatively similar synchronization crossovers, although the physics is different and in the quantum case a role in the crossover is played by entanglement.
The paper is organized as follows. Section 2 introduces the model in both the classical mean-field and quantum finite-size settings. Section 3 examines classical chaotic synchronization, focusing on Lyapunov exponents, Pearson coefficients, and magnetization regimes. Section 4 presents the finite-spin quantum dynamics, highlighting crossover behavior and the role of entanglement. Section 5 concludes with a discussion of the relation between classical and quantum chaotic synchronization.
2 Model
2.1 Lindblad equation and classical mean field limit
We consider a system consisting of two coherently coupled spins of magnitude $S$, evolving under a Lindblad dynamics as introduced in Ref. [39]. The unitary part of the evolution is provided by the following Hamiltonian (we will work in natural units, where $\hbar=1$),
$\hat{H}=\Omega\hat{S}_{A}^{x}+\Omega\hat{S}_{B}^{x}+\frac{\Gamma}{2S}\left(\hat{S}_{A}^{+}\hat{S}_{B}^{-}+\hat{S}_{A}^{-}\hat{S}_{B}^{+}\right)\,$, (1)
where $\hat{S}_{\alpha}^{j}$ can be considered as the sum of NN spin-$1/2$ degrees of freedom, $\hat{S}_{\alpha}^{j}=\frac{1}{2}\sum_{l=1}^{N}\hat{\sigma}_{j\,l}^{\alpha}$ (with $\alpha\in\{x,y,z\}$ and $j\in\{A,B\}$). Moreover, $\hat{S}_{j}^{\pm}=\hat{S}_{j}^{x}\pm i\hat{S}_{j}^{y}$. By choosing $S=N/2$, the system is restricted to the Hilbert subspace with maximum total spin of the NN spin-$1/2$ degrees of freedom. Therefore, the limit $S\to\infty$ corresponds to the thermodynamic limit, as shown in Ref. [36]. The parameter $\Gamma$ determines the strength of the coherent coupling between the subsystems and is renormalized by the factor $1/S$ to ensure the correct, extensive scaling of the interaction term in the thermodynamic limit. The parameter $\Omega$ is the driving frequency of the two subsystems when uncoupled ($\Gamma=0$).
Additionally, the evolution includes a dissipative component described by the Lindblad equation
$\frac{\mathrm{d}\hat{\rho}_{t}}{\mathrm{d}t}=-i\left[\hat{H},\hat{\rho}_{t}\right]+\frac{\kappa}{2S}\Big(\hat{S}_{A}^{-}\,\hat{\rho}_{t}\,\hat{S}_{A}^{+}+\hat{S}_{B}^{+}\,\hat{\rho}_{t}\,\hat{S}_{B}^{-}-\tfrac{1}{2}\{\hat{S}_{A}^{+}\hat{S}_{A}^{-}+\hat{S}_{B}^{-}\hat{S}_{B}^{+},\hat{\rho}_{t}\}\Big)\,$, (2)
where $\kappa$ is the decay (excitation) rate for subsystem AA ($B$). This dynamics conserves the variables $({\hat{S}_{j}})^{2}=({\hat{S}_{j}^{x}})^{2}+({\hat{S}_{j}^{y}})^{2}+({\hat{S}_{j}^{z}})^{2}$, with $j=A,B$, and – as mentioned above – we choose $({\hat{S}_{j}})^{2}=S(S+1)=\frac{N}{2}\left(\frac{N}{2}+1\right)$ for both j=A,Bj=A,\,B. We fix $\kappa=1$ throughout.
For any generic operator $\hat{\mathcal{O}}$, we can write the Heisenberg-like evolution of its expectation value as [36]
$\frac{\mathrm{d}}{\mathrm{d}t}\braket{\hat{\mathcal{O}}}_{t}$ $=i\braket{[\hat{H},\hat{\mathcal{O}]}}_{t}$
$+\frac{\kappa}{2S}\braket{\left\{\left[\hat{S}_{A}^{+},\hat{\mathcal{O}}\right]\hat{S}_{A}^{-}+\hat{S}_{A}^{+}\left[\hat{\mathcal{O}},\hat{S}_{A}^{-}\right]+\left[\hat{S}_{B}^{-},\hat{\mathcal{O}}\right]\hat{S}_{B}^{+}+\hat{S}_{B}^{-}\left[\hat{\mathcal{O}},\hat{S}_{B}^{+}\right]\right\}}_{t}\,$, (3)
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\braket{\hat{S}_{A}^{x}}_{t}$ $=\frac{\Gamma}{S}\braket{\hat{S}_{A}^{z}\hat{S}_{B}^{y}}+\frac{\kappa}{2S}\left(\{\hat{S}_{A}^{z},\hat{S}_{A}^{x}\}-\hat{S}_{A}^{x}\right)$ (4)
$\frac{\mathrm{d}}{\mathrm{d}t}\braket{\hat{S}_{B}^{x}}_{t}$ $=\frac{\Gamma}{S}\braket{\hat{S}_{B}^{z}\hat{S}_{A}^{y}}-\frac{\kappa}{2S}\left(\{\hat{S}_{B}^{z},\hat{S}_{B}^{x}\}-\hat{S}_{B}^{x}\right)$
$\frac{\mathrm{d}}{\mathrm{d}t}\braket{\hat{S}_{A}^{y}}_{t}$ $=-\Omega\braket{\hat{S}_{A}^{z}}_{t}-\frac{\Gamma}{S}\braket{\hat{S}_{A}^{z}\hat{S}_{B}^{x}}+\frac{\kappa}{2S}\left(\{\hat{S}_{A}^{z},\hat{S}_{A}^{y}\}-\hat{S}_{A}^{y}\right)$
$\frac{\mathrm{d}}{\mathrm{d}t}\braket{\hat{S}_{B}^{y}}_{t}$ $=-\Omega\braket{\hat{S}_{B}^{z}}_{t}-\frac{\Gamma}{S}\braket{\hat{S}_{B}^{z}\hat{S}_{A}^{x}}-\frac{\kappa}{2S}\left(\{\hat{S}_{B}^{z},\hat{S}_{B}^{y}\}-\hat{S}_{B}^{y}\right)$
$\frac{\mathrm{d}}{\mathrm{d}t}\braket{\hat{S}_{A}^{z}}_{t}$ $=\Omega\braket{\hat{S}_{A}^{y}}_{t}+\frac{\Gamma}{2S}\left(\hat{S}_{A}^{y}\hat{S}_{B}^{x}-\hat{S}_{A}^{x}\hat{S}_{B}^{y}\right)-\frac{\kappa}{S}\left[(\hat{S}_{A}^{x})^{2}+(\hat{S}_{A}^{y})^{2}\right]$
$\frac{\mathrm{d}}{\mathrm{d}t}\braket{\hat{S}_{B}^{z}}_{t}$ $=\Omega\braket{\hat{S}_{B}^{y}}_{t}+\frac{\Gamma}{2S}\left(\hat{S}_{B}^{y}\hat{S}_{A}^{x}-\hat{S}_{B}^{x}\hat{S}_{A}^{y}\right)+\frac{\kappa}{S}\left[(\hat{S}_{B}^{x})^{2}+(\hat{S}_{B}^{y})^{2}\right]\,$,
If we consider the reduced variables $\hat{m}_{A/B}^{\alpha}=\hat{S}_{A/B}^{\alpha}/S$ we see that they commute in the limit $S\to\infty$. In this limit, therefore, one can neglect quantum correlations for these variables. So, rewriting Eq. (4) for the reduced variables, taking the limit $S\to\infty$ and neglecting quantum correlations, one obtains
$\frac{\mathrm{d}m_{A}^{x}}{\mathrm{d}t}$ $=\kappa\,m_{A}^{x}m_{A}^{z}+\Gamma\,m_{A}^{z}m_{B}^{y}$ (5)
$\frac{\mathrm{d}m_{B}^{x}}{\mathrm{d}t}$ $=-\kappa\,m_{B}^{x}m_{B}^{z}+\Gamma\,m_{B}^{z}m_{A}^{y}$
$\frac{\mathrm{d}m_{A}^{y}}{\mathrm{d}t}$ $=-\Omega\,m_{A}^{z}+\kappa\,m_{A}^{y}m_{A}^{z}-\Gamma\,m_{A}^{z}m_{B}^{x}$
$\frac{\mathrm{d}m_{B}^{y}}{\mathrm{d}t}$ $=-\Omega\,m_{B}^{z}-\kappa\,m_{B}^{y}m_{B}^{z}-\Gamma\,m_{B}^{z}m_{A}^{x}$
$\frac{\mathrm{d}m_{A}^{z}}{\mathrm{d}t}$ $=\Omega\,m_{A}^{y}-\kappa\left((m_{A}^{x})^{2}+(m_{A}^{y})^{2}\right)+\Gamma\left(m_{A}^{y}m_{B}^{x}-m_{A}^{x}m_{B}^{y}\right)$
$\frac{\mathrm{d}m_{B}^{z}}{\mathrm{d}t}$ $=\Omega\,m_{B}^{y}+\kappa\left((m_{B}^{x})^{2}+(m_{B}^{y})^{2}\right)+\Gamma\left(m_{B}^{y}m_{A}^{x}-m_{B}^{x}m_{A}^{y}\right).$
2.2 Quantum-trajectory approach for the finite-size case
Let us consider Eq. (2) where, as stated above, we restrict to the tensor product subspace such that $({\hat{S}_{A}})^{2}=({\hat{S}_{B}})^{2}=S(S+1)$. Since the dynamics conserves both $({\hat{S}_{A}})^{2}$ and $({\hat{S}_{B}})^{2}$, we can fix these values throughout the evolution. Therefore we can notice that
$\hat{S}_{A}^{+}\hat{S}_{A}^{-}$ $=-({\hat{S}_{A}^{z}})^{2}+({\hat{S}_{A}})^{2}+\hat{S}_{A}^{z}=S\left(S+1\right)+\hat{S}_{A}^{z}-({\hat{S}_{A}^{z}})^{2}$ (6)
$\hat{S}_{B}^{-}\hat{S}_{B}^{+}$ $=S\left(S+1\right)-\hat{S}_{B}^{z}-({\hat{S}_{B}^{z}})^{2}\,.$
We implement this dynamics using the quantum-jump unraveling of the Lindblad equation Eq. (2). One can show that the Lindblad evolution is equivalent to the average over infinite realizations of a pure-state dynamics with a stochastic non-Hermitian Schrödinger equation. For each Lindblad equations there are many possible stochastic-dynamics approaches leading to it and are called unravelings (see Refs. [42, 43, 44] for a review). Here we choose the so-called quantum-jump unraveling that takes the following form. We discretize the time dividing it in intervals $\delta t$ and in each time interval:
  • With probability
    $p_{1}=\frac{\kappa}{2S}\delta t\braket{\psi_{t}|\hat{S}_{A}^{+}\hat{S}_{A}^{-}|\psi_{t}}=\frac{\kappa}{2S}\delta t\left[S\left(S+1\right)+\braket{\psi_{t}|[\hat{S}_{A}^{z}-(\hat{S}_{A}^{z})^{2}]|\psi_{t}}\right]$ (7)
    we perform the transformation
    $\ket{\psi_{t}}\to\ket{\psi_{t+\delta t}}=\frac{\hat{S}_{A}^{-}\ket{\psi_{t}}}{||\hat{S}_{A}^{-}\ket{\psi_{t}}||}\,.$ (8)
  • With probability
    $p_{2}=\frac{\kappa}{2S}\delta t\braket{\psi_{t}|\hat{S}_{B}^{-}\hat{S}_{B}^{+}|\psi_{t}}=\frac{\kappa}{2S}\delta t\left[S\left(S+1\right)-\braket{\psi_{t}|[\hat{S}_{B}^{z}+(\hat{S}_{B}^{z})^{2}]|\psi_{t}}\right]$ (9)
    we perform the transformation
    $\ket{\psi_{t}}\to\ket{\psi_{t+\delta t}}=\frac{\hat{S}_{B}^{+}\ket{\psi_{t}}}{||\hat{S}_{B}^{+}\ket{\psi_{t}}||}\,.$ (10)
  • With probability $p=1-p_{1}-p_{2}$ we evolve along $\delta t$ with the non-Hermitian Hamiltonian
    $\hat{H}_{\rm nH}=\Omega(\hat{S}_{A}^{x}+\hat{S}_{B}^{x})+\frac{\Gamma}{2S}\left(\hat{S}_{A}^{+}\hat{S}_{B}^{-}+\hat{S}_{A}^{-}\hat{S}_{B}^{+}\right)-i\frac{\kappa}{4S}\left(\hat{S}_{A}^{+}\hat{S}_{A}^{-}+\hat{S}_{B}^{-}\hat{S}_{B}^{+}\right)\,$, (11)
    or equivalently with
    $\hat{H}_{\rm nH}^{\prime}=\Omega(\hat{S}_{A}^{x}+\hat{S}_{B}^{x})+\frac{\Gamma}{2S}\left(\hat{S}_{A}^{+}\hat{S}_{B}^{-}+\hat{S}_{A}^{-}\hat{S}_{B}^{+}\right)-i\frac{\kappa}{4S}\left(\hat{S}_{A}^{z}-\hat{S}_{B}^{z}-(\hat{S}_{A}^{z})^{2}-(\hat{S}_{B}^{z})^{2}\right)\,.$ (12)
    (The two Hamiltonians are equal up to a constant.) Writing $\hat{H}_{\rm nH}^{\prime}=\hat{H}-i\hat{K}$, we evolve by trotterization
    $\ket{\psi_{t}}\to\ket{\psi_{t+\delta t}}=\frac{\operatorname{e}^{-i\hat{H}\delta t}\operatorname{e}^{\hat{K}\delta t}\ket{\psi_{t}}}{||\operatorname{e}^{-i\hat{H}\delta t}\operatorname{e}^{\hat{K}\delta t}\ket{\psi_{t}}||}\,.$ (13)
We initialize the system with both collective spins pointing upwards, $S_{A}^{z}=S_{B}^{z}=S$ and $S_{A}=S_{B}=S$, and fix everywhere $\kappa=1$.
3 Mean-field dynamics
3.1 Methods
3.1.1 Equations.
[…]
3.1.2 Initialization.
Initial conditions are generated by taking $(m_{j}^{x},m_{j}^{y},m_{j}^{z})=(e,\,d,\,\sqrt{1-e^{2}-d^{2}})$ with ($e$, $d$) uniformly sampled within a disk of radius $a$. The sampling satisfies the constraint
$e^{2}+d^{2}\leq a^{2}.$ (14)
In polar coordinates we can write for all $j$ $(m^{x},m^{y},m^{z})=(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta)$ and $0<\sin\theta<a$, that is to say $0<\theta<\arcsin a$. The variables $e$ and $d$ are jointly uniformly distributed over a disk of radius $a$, meaning their joint probability density function is
$p(e,d)=\frac{1}{\pi a^{2}}\,.$ (15)
Therefore we average over a cap surrounding the north pole of the sphere. Notice that $a=1$ corresponds to averaging over the whole northern hemisphere.
3.1.3 Largest Lyapunov exponent.
To evaluate the chaoticity of the dynamics we evaluate the largest Lyapunov exponent $\Lambda_{L}$ [45] (LLE), defined as the limit of the logarithmic rate of separation of two initially close trajectories
$\Lambda_{L}=\lim_{d_{0}\to 0}\lim_{t\to\infty}\frac{1}{t}\ln\left[\frac{d(t)}{d_{0}}\right]\,$, (16)
where $d_{0}$ is the initial distance between the two trajectories, and $d(t)$ is the distance at time $t$. To compute the Lyapunov exponent numerically, we use the method described in Ref. [46]. We evolve two initially close trajectories according to Eq. (5). We discretize the time with steps $\delta t$ and at each time step $n$, we calculate the distance $d_{n}$ between the two trajectories, and evaluate the logarithm of the ratio of $d_{n}$ and $d_{0}$. Then we move one of the phase points along the line joining the two points in order to restore the initial distance, and repeat the process. The set of obtained logarithms is then averaged over time to obtain the Lyapunov exponent
$\Lambda_{L}=\frac{\delta t}{T_{\rm max}}\sum_{n=1}^{T_{\rm max}/\delta t}\ln\left(\frac{d_{n}}{d_{0}}\right)\,.$
Specifically, we perform the averaging over a time interval of length $T_{\text{max}}$, call the number of steps used as $K\equiv T_{\rm max}/\delta t$, and average $\Lambda_{L}$ over a set of randomly chosen initial conditions, as specified in Sec. 3.1.2. We mark the Lyapunov exponent averaged over random initial conditions as $\braket{\Lambda_{L}}$.
3.1.4 Pearson coefficient.
To study the correlations between the subsystems, we calculate the Pearson coefficient $C_{P}$ (see for instance Refs. [18, 47]). To compute this coefficient for the time evolution of subsystems $A$ and $B$, we use the following approach. We take a set of random initial conditions as specified in Sec. 3.1.2. At each time $t$ we evaluate the Pearson correlation coefficient between different subsystems as
$C_{P}(t)=\frac{\braket{m_{A}^{z}(t)m_{B}^{z}(t)}-\braket{m_{A}^{z}(t)}\braket{m_{B}^{z}(t)}}{\sqrt{\braket{(m_{A}^{z})^{2}(t)}-\braket{m_{A}^{z}(t)}^{2}}\sqrt{\braket{(m_{B}^{z})^{2}(t)}-\braket{m_{B}^{z}(t)}^{2}}}\,$, (17)
[…]
3.2 Results
Figure 1: Largest Lyapunov exponent $\Lambda_{L}$ versus $\Gamma$ and $\Omega$ for initial conditions $m_{j}^{z}=1$ and $m_{j}^{x,y}=0$ for all $j$ (panel a), averaged over random initial conditions as given in Sec. 3.1.2 with $a=0.1$ (panel b), average over random initial conditions as given in Sec. 3.1.2 with $a=1$ (panel c). The averages are over $N_{\rm r}=50$ initial conditions. See the main text and Fig. 2 for a discussion about the symbols shown in panel (b).
Figure 2: Largest Lyapunov exponent $\braket{\Lambda_{L}}$ (top row) and Pearson correlation coefficient $\overline{C_{P}}$ (bottom row) versus $\Gamma$ along the line $\Gamma+\Omega=1.5$ (left column), $\Gamma+\Omega/2=2$ (center column), and $\Gamma+\Omega/2=1.5$ (right column). These values are calculated as averages over $N_{\rm r}=100$ initial conditions generated on a spherical cap with parameter $a=0.1$.
We begin by plotting the LLE defined in Sec. 3.1.3, shown in Fig. 1. In panel (a) we choose a specific initial condition (all spins pointing along the north pole), while in panel (b) and (c) we average over initial conditions as given in Sec. 3.1.2, choosing respectively $a=0.1$ and $a=1$. All the regimes described in Ref. [39] for Fig. 1(a) are clearly visible also when we average over many initial conditions, though they appear smeared out. In particular, at the bottom left there is a regime of negative LLE (absence of chaos) corresponding to relaxation of the two magnetizations to a fixed point (trivial phase). Then there is a central lobe where the LLE is vanishing that corresponds to a continuous time crystal phase (marked as CTC3 in Ref. [39]). Another similar lobe appears in the upper left region, where there is another continuous time crystal phase (CTC1) which encompasses the usual dissipative time crystal (occurring when the coupling $\Gamma$ vanishes).
Figure 3: Heat maps of various quantities versus $\Gamma$ and $\Omega$. (a,b) Pearson correlation coefficient. (c,d) Difference between the magnetizations averaged over time and initial conditions, $\langle\overline{\delta m}\rangle=\braket{\overline{m_{B}^{z}}}-\braket{\overline{m_{A}^{z}}}$. (e,f) Relative sign $s=\operatorname{sign}(\braket{\overline{m_{z}^{A}}}\braket{\overline{m_{z}^{B}}})$ of the two magnetizations. The calculations are performed by averaging over $N_{\rm r}=100$ random initial conditions generated on a spherical cap with parameter $a=0.1$ (first row), and $a=1$ (second row).
In Fig. 1(b), neglecting the trivial phase, we consider three lines cutting the parameter space in Fig. 1(b) from the upper left to the bottom right part of the panel; These lines are $\Omega+\Gamma=1.5$, $\Gamma+\Omega/2=2$, and $\Gamma+\Omega/2=1.5$ and are marked in Fig. 1(b) with different symbols (blue circles, green triangles, and orange squares, respectively). Along these lines, we plot, using the same symbols, the LLE (Fig. 2, top row) and the corresponding average Pearson coefficient described in Sec. 3.1.4 (Fig. 2, bottom row) versus $\Gamma$. Both the LLE and the Pearson coefficient are evaluated over a set of random initial conditions with $a=0.1$. We see that there is a threshold value $\Gamma\simeq 1.2$ in panel 2(a), $\Gamma\simeq 1.6$ in panel 2(b), $\Gamma\simeq 1.3$ in panel 2(c)] where the LLE moves from zero to positive, coinciding with a sudden increase of the Pearson coefficient. At this threshold, the Pearson coefficient shifts from strongly negative to strongly positive values. Below this threshold the system is in the CTC3 phase: Here, the Lyapunov exponent is vanishing, and the negative Pearson coefficient indicates that the two subsystem oscillate out of phase with respect to each other. Above the threshold the system becomes chaotic – as marked by the Lyapunov exponent becoming positive – but the two oscillators behave in a strictly correlated way, as witnessed by the positive value of the Pearson coefficient. Therefore, in this regime, the two subsystems display full chaotic synchronization.
The full chaotic synchronization regime is better highlighted in Fig. 3(a,b), showing heat maps of the Pearson coefficient across the parameter space. In Fig. 3(a) we show the heat map of the Pearson coefficient averaged over a cap with $a=0.1$ and we can see a close correspondence of this plot with the heat map of the LLE evaluated with the same initial conditions [see Fig. 1(b)].
The behaviors of the Lyapunov exponent and Pearson coefficient closely relate to the behavior of the magnetization averaged over time and initial-condition $m_{j}^{z}(t)$ ($j=A,B$). We define this average as $\braket{\overline{m_{j}^{z}}}$ and in Fig. 4 we plot it versus $\Gamma$ along the three lines considered above. We identify three regimes: (i) a small-$\Gamma$ regime, where the average magnetization is vanishing for both subsystems; (ii) an intermediate-$\Gamma$ regime, where the average magnetizations of the two subsystems have opposite sign (staggered magnetization regime); (iii) a large-$\Gamma$ regime, where the average magnetizations of the two subsystems are nonvanishing and nearly equal (uniform magnetization regime). We find that the crossover between the staggered and uniform magnetization regimes coincides with the onset of full chaotic synchronization, as we can see by comparing the top row of Fig. 4 (where initial conditions are with $a=0.1$) with Fig. 2.
Figure 4: Time and initial-condition average values of $m_{j}^{z}(t)$, $\braket{\overline{m_{j}^{z}}}$ versus $\Gamma$, along the line $\Omega+\Gamma=1.5$ (a, left column), $\Gamma+\Omega/2=2$ (b, center column), and $\Gamma+\Omega/2=1.5$ (c, right column). We average over a set of $N_{\rm r}=100$ initial conditions generated either for $a=0.1$ (top row) or $a=1$ (bottom row). The legend is shared.
In order to better highlight these regimes, we plot in Fig. 3(c,d) the heat maps for the difference between the magnetizations averaged over time and initial conditions, $\delta m\equiv\braket{\overline{m_{B}^{z}}}-\braket{\overline{m_{A}^{z}}}$ [in panel (c) we take $a=0.1$, in panel (d) $a=1$], and compare them with the heat maps of the Pearson coefficient, with the corresponding value of $a$ [Fig. 3(a,b)]. We see two regions where $\langle\overline{\delta m}\rangle\approx 0$, on the top left and the bottom right, both corresponding to positive values of the Pearson coefficient. The top-left region corresponds to regime (i) with vanishing magnetization, while the bottom-right region corresponds to regime (iii), characterized by uniform magnetization. We see also a central lobe where the difference $\delta m$ is slightly positive and the Pearson coefficient is negative. This region corresponds to the staggered magnetization regime (ii). We also plot the heat map of the relative sign of the two averaged magnetizations, i. e., $s=\operatorname{sign}(\braket{\overline{m_{A}^{z}}}\braket{\overline{m_{B}^{z}}})$ [Fig. 3(e,f)], and see a clear positive region at the bottom right. This is the region where regime (iii) occurs and – as we can see comparing with Fig. 1(b,c) and Fig. 3(a,b) – it corresponds to a positive Lyapunov exponent and to a positive Pearson coefficient that suddenly increases at the boundary of this region, as we have described above. Thus, this is a region of chaotic synchronization, that is therefore closely related to the presence of a uniform magnetization.
4 Quantum finite-size dynamics
Figure 5: (a,c) $\overline{S_{A}^{z}}$ and $\overline{S_{B}^{z}}$ versus $\Omega$ for different NN. (b,d) Time-averaged connected correlator $\overline{C_{AB}^{zz}}$. Average over $N_{\rm r}\geq 48$ trajectories. On the left column we consider the curves for the line $\Omega/2+\Gamma=2$ and on the right one $\Omega/2+\Gamma=1.5$.
[…]
$\overline{C_{AB}^{zz}}=\overline{\braket{S_{A}^{z}S_{B}^{z}}}-\overline{\braket{S_{A}^{z}}\braket{S_{B}^{z}}}.$ (19)
Further insight is gained by analyzing histograms of these quantities over time $t$ and across different trajectory realizations $r$ (as done in Refs. [48, 49]). We plot the histograms for $(S_{A}^{z})^{r}(t)=\braket{\psi_{t}^{r}|\hat{S}_{A}^{z}|\psi_{t}^{r}}$ and $(S_{B}^{z})^{r}(t)=\braket{\psi_{t}^{r}|\hat{S}_{B}^{z}|\psi_{t}^{r}}$ in Fig. 6(a,b). They are obtained as an histogram over a set of values obtained by changing $r$ and $t$. Along each trajectory, we also consider the entanglement entropy between the two subsystems, defined as
[…]
and plot its histograms in Fig. 6(c). All histograms have a maximum, and we plot its position as a function of $\Gamma$ in Fig. 7. We can clearly observe the crossover from the staggered to the uniform magnetized regime in the behavior of $(S_{A}^{z})^{\rm max}$ and $(S_{B}^{z})^{\rm max}$ [Fig. 7(a,c)], also accompanied by a peak in the behavior of the maximum point of the entanglement entropy $\mathcal{E}^{\text{max}}$ [Fig. 7(b,d)].
Figure 6: (a) Histogram of $(S_{A}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (b) Histogram of $(S_{B}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (c) Histogram of the entanglement entropy $\mathcal{E}^{r}(t)$ as it changes along time $t$ and across trajectories $r$. We fix the constraint $\Omega/2+\Gamma=1.5$, $\delta t=10^{-4}$ and $S=20$.
Because the classical system’s crossover closely relates to full chaotic synchronization, we interpret these quantum observations as evidence of quantum synchronization. In analogy with the classical mean-field case, we can say that the quantum synchronization sets in when the magnetization suddenly becomes uniform. Moreover, as in the classical case, also in the quantum one the synchronization is chaotic, as one can argue considering the properties of the average density matrix
$|\hat{\rho}_{t}=\frac{1}{N_{\rm r}}\sum_{r=1}^{N_{\rm r}}\ket{\psi_{t}^{r}}\bra{\psi_{t}^{t}}$ (21)
for a time long enough that the steady state has been reached. This is the so-called nonequilibrium steady state (NESS) density matrix, $\hat{\rho}_{\rm NESS}$. After obtaining $\hat{\rho}_{\rm NESS}$ as an average over realizations at a time long enough that all finite-time transients have died out, we evaluate the average level spacing ratio $r_{S}$ of this matrix as described in Ref. [50]. Denoting $\rho_{\alpha}$ the eigenvalues of $\hat{\rho}_{\rm NESS}$ taken in increasing order, the average level spacing ratio is defined as
$r_{S}=\frac{1}{2S-1}\sum_{\alpha=1}^{2S-1}\frac{\min(\rho_{\alpha+1}-\rho_{\alpha},\rho_{\alpha+2}-\rho_{\alpha+1})}{\max(\rho_{\alpha+1}-\rho_{\alpha},\rho_{\alpha+2}-\rho_{\alpha+1})}\,.$ (22)
Evaluating it, we always find $r_{s}\simeq 0.59$ (see Fig. 8), a value only slightly smaller than the random-matrix Gaussian unitary ensemble value $r_{\rm GUE}\simeq 0.5996$ resulting from a numerical fit in Ref. [51]. We can conclude that the steady-state Lindblad density matrix behaves as a random matrix, a behavior that corresponds to quantum chaos [50]. Beyond that, it displays a sudden crossover from staggered to uniform magnetization strictly reminiscent of the synchronization phenomenology observed in the mean field [see Fig. 7(a,c)]. Therefore, also in the quantum finite-$S$ case, we can identify this phenomenon as quantum chaotic synchronization.
Figure 7: (a,c) Maximum points of the histograms of $S_{A}^{z}$ and $S_{B}^{z}$ along and across trajectories versus $\Gamma$. (b,d) Maximum point of the histogram of the entanglement entropy versus $\Gamma$. We fix the constraint $\Omega/2+\Gamma=2$ in the left column and $\Omega/2+\Gamma=1.5$ in the right column. We take also $\delta t=10^{-4}$.
We notice that the threshold for this crossover in the quantum case is smaller than in the mean field case [compare Fig. 7(a,c) with the corresponding Fig. 2(b,c)]. The point is that the infinite-$S$ limit and the infinite time-limit do not commute. Indeed, the infinite-$S$ limit shows persistent (periodic or aperiodic) oscillations, while the finite-$S$ size case relaxes for $t\to\infty$ to a NESS density matrix for any arbitrary large but finite $S$ [39]. Here we are studying the properties of this NESS density matrix, which are in general different from the ones of the long-time behavior in the mean-field regime (see examples in Refs. [52, 41]). Thus, it is not strange that the crossover occurs at different values of $\Gamma$ for the mean-field infinite-$S$ regime and for the long-time limit of the quantum finite-$S$ case. What is remarkable is that the crossover is present in both scenarios, despite the non-commutativity of the limits and their different long-time behaviors; additionally, entanglement plays a significant role in the quantum case but is absent classically. We also note that the noncommutative nature of these limits helps explain why the largest Lyapunov exponent (LLE), shown in Fig. 2, exhibits intervals of zero (regular dynamics) and positive (chaotic dynamics), whereas the NESS density matrix displays signatures of quantum chaos throughout the same parameter range (see Fig. 8). This difference arises from the LLE assessing the long-time dynamics of the mean-field, infinite-$S$ limit, whereas the NESS characterizes the long-time limit of finite-size quantum dynamics – the infinite-$S$ limit being taken first in the former case.
Figure 8: Average level spacing ratio of the NESS density matrix versus $\Gamma$ for different values of $S$, compared with the average level spacing ratio for the Gaussian Unitary Ensemble as evaluated in Ref. [51]. The flatness of the curve and the agreement with the GUE value increase for increasing subsystem spin magnitude $S$. We fix the constraint $\Omega/2+\Gamma=2$ and evaluate $\hat{\rho}_{\rm NESS}$ as $\hat{\rho}_{t}$ with $t=100$ in units of $1/\kappa$. In Eq. (21) we average over $N_{\rm r}=48$ quantum trajectories and take $\delta t=10^{-4}$.
5 Conclusions
In conclusion, we have investigated the dynamics of two coupled dissipative time crystals, identifying signatures of chaotic synchronization in both the classical and quantum regimes. In the mean-field infinite-$S$ limit, we found a region of parameter space characterized by a positive largest Lyapunov exponent and a Pearson correlation coefficient close to one, signaling the coexistence of chaos and strong synchronization [30, 31, 28]. At the onset of this regime, the Pearson coefficient increases abruptly, in correspondence with a transition in the structure of the time-averaged magnetizations. Specifically, chaotic synchronization is associated with a uniform nonvanishing magnetization across the subsystems, whereas the staggered-magnetization regime corresponds to a regular-dynamics time-crystal phases (CTC3 phase of Ref. [39]), and the vanishing-magnetization one encompasses to a chaotic unsynchronized phase and another (CTC1) time crystal phase. Both the CTC1 and CTC3 phases display a nonvanishing Pearson coefficient (and then synchronization), but in those phases the Lyapunov exponent is vanishing and there is no chaos. The presence of both chaos and nonvanishing Pearson coefficient appears only in the uniform-magnetization regime, and this establishes a direct correspondence between chaotic synchronization and collective order.
We then turned to finite-$S$ systems, where quantum fluctuations become relevant. Using a quantum-trajectory approach, we analyzed subsystem magnetizations along stochastic trajectories and their distributions over time and trajectory realizations. The maxima of these histograms exhibit a sharp crossover between staggered and uniform magnetization, closely resembling the mean-field case and indicating a quantum analogue of chaotic synchronization. Moreover, dissipative quantum chaos is revealed by Gaussian unitary ensemble statistics in the nonequilibrium steady-state density matrix [50]. Beyond that, the entanglement entropy between subsystems displays a correlated feature at the crossover, showing that entanglement contributes to the quantum synchronization process [13, 4, 5].
In summary, this model displays classical chaotic synchronization and an analogous quantum phenomenon, which we can call quantum chaotic synchronization. Interestingly, the crossover between the staggered and the uniform magnetization in the mean-field classical chaotic synchronization occurs at a different point in parameter space than the finite-$S$ threshold between staggered and uniform magnetization in the quantum case. This difference arises because the limits $S\to\infty$ and $t\to\infty$ do not commute in this quantum dynamics. In the mean-field case, the $S\to\infty$ limit is taken first, resulting in persistent oscillations that can be periodic or aperiodic [39]. In the quantum case, the $t\to\infty$ limit comes first and the system always relaxes to a NESS. Remarkably, the NESS shares qualitative features with the mean-field state, despite their fundamental differences and the presence of entanglement only in the quantum case. The noncommutativity of these limits further explains the discrepancies in parameter regimes where classical and quantum chaos appear.
Our findings provide a consistent picture of chaotic synchronization between dissipative time crystals, in both the classical and the quantum regime, and support the identification of a phenomenon of quantum chaotic synchronization. Future research developments include the extension of this analysis to the case of a chain of many coupled subsystems (possibly with long-range dissipation [53]), the study of classical and quantum correlation in the case of classical and quantum coupled dissipative kicked rotors [54], and the exploration of the relationship between chaotic synchronization and other kinds of quantum resources [55].
We acknowledge P. Solanki for fruitful discussions. G. P. and A. R. acknowledge financial support from PNRR MUR Project No. PE0000023-NQSTI. We acknowledge computational resources from the CINECA award under the ISCRA initiative and from MUR, PON “Ricerca e Innovazione 2014-2020,” under Grant No. PIR01_\_00011 - (I.Bi.S.Co.). This work was supported by PNRR MUR Project No. PE0000023-NQSTI, by the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733, by MUR Project No. CN_\_00000013-ICSC (P. L.), and by the QuantERA II Programme STAQS project, which received funding from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733 (P. L.).